OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. sin x/ 1 - cos x + sin x/ 1 + cos x = 2 csc x

OpenStudy (freckles):

sin(x)/(1-cos(x))+sin(x)/(1+cos(x)) for the left hand side ? If so I think it should first be thought to combine the fractions since our goal is to right the left hand side as one term (we do after all have one term on the right)

OpenStudy (freckles):

to write the left*

OpenStudy (anonymous):

What would I do to combine?

OpenStudy (freckles):

find a common denominator

OpenStudy (freckles):

or force once

OpenStudy (freckles):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+cb}{bd} \text{ \to force a common denominator }\]

OpenStudy (anonymous):

ahh I see! Give me one moment to do that

OpenStudy (anonymous):

\[sinx(1 + cosx) + sinx(1 - cosx)/ (1 - cosx)(1 + cosx)\]

OpenStudy (freckles):

multiply the top out and the bottom out something should cancel on top and you should be able to use a really cool identity on bottom after the multiplying part

OpenStudy (anonymous):

What would I do from here?

OpenStudy (anonymous):

sorry I lost connection...

OpenStudy (freckles):

Did you try to multiply the top and bottom out?

OpenStudy (anonymous):

For the botton I got 2-cos^2x

OpenStudy (anonymous):

I'm not sure what to do with the top

OpenStudy (freckles):

\[\frac{\sin(x)+\sin(x) \cos(x)+\sin(x)-\sin(x)\cos(x)}{1-\cos^2(x)}\] the bottom should be 1-cos^2(x)

OpenStudy (anonymous):

oh ok

OpenStudy (freckles):

\[(1-\cos(x))(1+\cos(x)) \\ =1(1+\cos(x))-\cos(x)(1+\cos(x)) \\ =1+\cos(x)-\cos(x)-\cos^2(x) \\ =1+0-\cos^2(x) \\=1-\cos^2(x) \] so that is how I got the bottom

OpenStudy (freckles):

for the top I just multiplied as well

OpenStudy (freckles):

sin(x)(1+cos(x)) is sin(x)+sin(x)cos(x) and sin(x)(1-cos(x)) is sin(x)-sin(x)cos(x)

OpenStudy (freckles):

you should see something cancel on top

OpenStudy (freckles):

and you should also be able to combine any like terms on top

OpenStudy (freckles):

and 1-cos^2(x) should look familiar (hint pythagorean identity)

OpenStudy (anonymous):

hold on I finally understand it

OpenStudy (anonymous):

1-cos^2x = sin^2x

OpenStudy (freckles):

yep

OpenStudy (anonymous):

so I should now get \[2\sin^2x \div \sin x^2\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

or is it just 2sinx?

OpenStudy (freckles):

do you know that x+x=2x and not 2x^2?

OpenStudy (freckles):

like 2+2=4 not 2(2)^2

OpenStudy (anonymous):

ok I thought I had made that mistake

OpenStudy (freckles):

if you have 1 smiley and you add another smiley you say you have 2 smiley not 2 smiley square

OpenStudy (freckles):

\[\frac{2 \sin(x)}{\sin^2(x)}\]

OpenStudy (freckles):

do you know where to go from here?

OpenStudy (anonymous):

what would we do now?

OpenStudy (anonymous):

hmmm

OpenStudy (freckles):

do you not see a common factor on top and bottom

OpenStudy (anonymous):

yes, so simplify?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

2\[\frac{ 2 }{ sinx}\]

OpenStudy (freckles):

yeah

OpenStudy (freckles):

2csc(x)

OpenStudy (anonymous):

and since sin x = 1/csc x you end up getting 2csc(x) correct?

OpenStudy (freckles):

yea

OpenStudy (anonymous):

Thank you so much! I have been really struggling with this unit.